OpenStudy (anonymous):

A sparrow flies to see his friend at a speed of 4km/h. His friend is out, so the sparrow immediately returns home. The whole journey took 54mins. How far away does his friend live?

OpenStudy (hoblos):

4km/60min in 54min sparrow flies 4*54/60 = 3.6km 3.6km is the distance covered for going and coming back so his friend live 1.8km away

OpenStudy (anonymous):

distance travelled in 60 min = 4000 m distance travelled in 1 min = 4000m/60 distance travelled in 54 min = (4000 * 54)/60 distance travlled one way = (4000 * 54)/60*2 = 1800 m = 1.8 km

OpenStudy (anonymous):

my text gives the answer 2km

OpenStudy (anonymous):

so the friend lives 1.8 km away

OpenStudy (anonymous):

it is rounded off to the nearest km...

OpenStudy (anonymous):

check if it says somewhere that you hv to give answer to the nearest km